Find f'(x) of f(x)= 1/{x-[2/(x + sinx)]}


1
Expert's answer
2021-03-08T17:16:45-0500

"f(x)=1\/(x-2\/(x+sinx))"

"f`(x)=-1\/(x-2\/(x+sinx))^2*(x-2\/(x+sinx))`"

"(x-2\/(x+sinx))`=1+2\/(x+sinx)^2(x+sinx)`"

"(x+sinx)`=1+cosx"

"f`(x)=-1\/(x-2\/(x+sinx))^2*(1+2\/(x+sinx)^2*(1+cosx))=\n((x+sinx)^2 + 2 +2cosx)\/(x(x+sinx)-2)^2"